Write the following numbers at the product of its prime factors in exponential form.

1.190   2.900   3.1323


1
Expert's answer
2021-05-02T07:14:45-0400

1.


"190=2\cdot95=2\cdot5\cdot19"

"190=2^1\cdot5^1\cdot19^1"



2.


"900=2\cdot450=2^2\cdot225=2^2\cdot3\cdot75"

"=2^2\cdot3^2\cdot25=2^2\cdot3^2\cdot5^2"


"900=2^2\cdot3^2\cdot5^2"



3.


"1323=3\cdot441=3^2\cdot147=3^3\cdot49"

"=3^3\cdot7^2"

"1323=3^3\cdot7^2"